40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 219 of 355  Not logged in ELOG logo
ID Date Author Type Category Subject
  6852   Fri Jun 22 03:37:42 2012 KojiUpdateGreen Locking2 arm ALS - Success!!!!

Are these correct?

1. It is a nice work.

2. This is not locking, but stabilization of the both arms by ALS.

3. We now have the phase trackers for both arms.

4. There is no coarse (i.e. short) delay line any more.

5. The splitters after the PDs are reducing the RF power to Beat-box.
Actually there are RF monitors on Beat-box for this purpose, but you did not notice them.

6. c1ioo channel list
https://wiki-40m.ligo.caltech.edu/CDS/C1IOO%20channel%20list
has to be updated.

7. Video can be uploaded to Youtube as Mike did at http://nodus.ligo.caltech.edu:8080/40m/6513

  6851   Fri Jun 22 02:21:57 2012 JenneUpdateGreen Locking2 arm ALS - Success!!!!

[Yuta, Jenne]

We locked both arms using the ALS system simultaneously!  Hooray!

Video of spectrum analyzer during lock acquisition of both beats is attached.

Jamie is super awesome, since he fixed us up a beatbox speedy-quick.  Thanks Jamie!!  speedy_gonzales-5257.jpg

 Details:

1:  Aligned PSL green optics

     1.1:  We added an amplifier of ~20dB after the X beat PD (more Xgreen power on the PSL table so the signal was ~3dB higher than Y, so required less amplification).  The ~24dB amplifier is still in place after the Y beat PD.  Both beat signals go to a splitter after their amplifiers.  One side of each splitter goes to one of the channels on the beatbox.  The other side of each splitter goes to a 3rd splitter, which we're using backwards to combine the 2 signals so we can see both peaks on the spectrum analyzer at the same time.

2:  Found both beat notes

     2.1:  Y beat was easy since we knew the temps that have been working for the past several days

     2.2:  X beat was more tricky - the last time it was locked was the end of February (elog 6342)

         2.2.1:  We found it by adjusting the PSL laser temp nearly the full range - DC Adjust slider was at 8.8V or so (Y beat was found with the slider at ~1.1V tonight)

          2.2.2:  We then walked the beat around to get the PSL temp back to "normal" by moving the PSL temp, then compensating with the Xend laser temp, keeping the beatnote within the range of the spectrum analyzer.

          2.2.3:  Fine tuned the temps of all 3 lasers until we had 2 peaks on the analyzer at the same time!!

               2.2.3.1:  Yend - measured Temp=34.14 C, thermal Out of Slow servo=29820

               2.2.3.2:  Xend - displayed temp=39.33 C, thermal Out of Slow servo=5070

               2.2.3.3: PSL - displayed temp=31.49 C, Slow actuator Adjust=1.100V

3:  Locked both arms using ALS!!

     3.1:  We were a little concerned that the Xarm wasn't locking.  We tried switching the cables on the beatbox so that we used the old channels for the Xarm, since the old channels had been working for Y.  Eventually we discovered that the input of the filter module for ETMX's POS-ALS input was OFF, so we weren't really sending any signals to ETMX.  We reverted the cabling to how it was this evening when Jamie reinstalled the beatbox.

          3.1.1:  We need to sort out our SUS screens - Not all buttons in medm-land link to the same versions of the SUS screens!  It looks like the ALS screen was modified to point the ETMY button to a custom ETMY SUS screen which has the ALS path in the POS screen, along with LSC and SUSPOS.  There is no such screen (that I have found) for ETMX.  The regular IFO_ALIGN screen points to the generic SUS screens for both ETMY and ETMX, so we didn't know until Yuta searched around for the filter bank that the ALS input for ETMX was off.  We just need to make sure that all of the screens reflect what's going on in the models.

     3.2:  See the video attached - it shows the beat peaks during locking!!! (how do I embed it? right now you have to download it)

          3.2.1:  First you will see both peaks moving around freely

          3.2.2:  Then X arm is locked briefly, then unlocked

          3.2.3: Y arm is locked, steadily increasing gain

          3.2.4:  X arm is locked, so both arms locked simultaneously

          3.2.5:  Yuta clicked a button, accidentally unlocking the Xarm

4:  The transmission of the X arm was not so great, and both of our green beams (although X green especially) were no longer nicely aligned with the cavities.  Yuta tried to align the X arm to the X green, but it's bad enough that we really need to start over with the whole IFO alignment - we leave this until tomorrow.  Since we didn't have any good IR transmission, we didn't bother to try to find and hold the Xarm on IR resonance using ALS, so we didn't measure a POX out of loop residual cavity motion spectrum.  Again, tomorrow. 

Attachment 1: P6210140.AVI
  6850   Thu Jun 21 20:07:18 2012 JamieUpdateGreen LockingImproved beatbox returns

I've reinstalled the beatbox in the 1X2 rack.  This improved version has the X and Y arm channels stuffed, but just one of the DFD channels (fine) each.

I hooked up the beat PD signals for X and Y to the RF inputs, and used the following two delay lines:

  • X: 140' light-colored cable on spool
  • Y: 30m black cable

The following channel --> c1ioo ADC --> c1gcv model connections were made:

  • X I  --> SR560 whitening --> ADC 22 -->  X fine I
  • X Q --> SR560 whitening --> ADC 23 --> X fine Q
  • Y I --> SR560 whitening --> ADC 24 --> Y fine I
  • Y Q --> SR560 whitening --> ADC 25 --> Y fine Q

The connections to the course inputs on the ALS block were grounded.  I then recompiled, reinstalled, and restarted c1gcv.  Functioning fine so far.

 

  6849   Thu Jun 21 15:36:51 2012 yutaUpdateLockingX arm alignment

I aligned X arm so that the beam spot comes roughly on the center.

1. Use ITMX and ETMX (mainly ITMX) to make beam spot come on center of the optic using eyeball.

2. Use ETMX and BS to maximize TRX power (reached ~ 0.85)

3. Aligned green optics on X end. Transmission of X green measured at PSL table is now 255 uW and TEM00 has the most power.

It was not easy to increase X green transmission more because beam spot on green transmission PD is wiggly when X end table is opened. When closed, wiggliness is about the same for Y green and X green.
Turning off HEPA on the X end didin't helped, but there must be something bad in the X end table. If we couldn't figure out why, let's wait for PZTs to come for end tables.

Considering the laser power is different(X end 1 W, Y end 700 mW), X green transmission should reach ~400 uW. But I think we should go on to X beat search.

I placed green shutter for X end back for convenience. I put some spacers to adjust its height and avoid beam clipping.


[Steve, Yuta]

What causing wiggly X green transmission was the air flow from the air conditioner. When we turned it off, beam spot motion became quiet. Air flow from HEPA was not effecting much.

  6848   Thu Jun 21 15:00:55 2012 steveUpdateSAFETYSURFs 2012 safety training

Masha, Eric, Yaakov, Liz and Sasha received 40m specific basic safety training.

Attachment 1: surfs2012.JPG
surfs2012.JPG
  6847   Thu Jun 21 12:56:49 2012 yutaUpdateLockingETMX 1064 trans camera

Quote:

[Jenne, Yuta]

We made ETMXT camera working.

 Xarm_EndTableLayout_NewTransCamera.png

Here's the new end table layout, for the transmitted IR stuff.

  6846   Thu Jun 21 12:13:35 2012 JenneUpdatePEMflow bench must be running all times

Quote:

The south end flow bench HEPA filter should be run all times. You can turn it off for a measurement or two but remember we are storing clean optics there.

The zero count bench will reach  room particle count  ~ 10,000 in one minute.

 My bad.  I turned it off last night to see if it would help make the Xgreen more stable, and then when I woke up this morning I realized that I had forgotten to turn it back on.  Bad Jenne.

  6845   Thu Jun 21 09:10:08 2012 steveUpdatePEMflow bench must be running all times

The south end flow bench HEPA filter should be run all times. You can turn it off for a measurement or two but remember we are storing clean optics there.

The zero count bench will reach  room particle count  ~ 10,000 in one minute.

  6844   Thu Jun 21 09:01:18 2012 steveUpdateSTACISthe resurrection of STACIS -surf

There is a small wood cabinet under the south end flow bench, labeled STACIS.

Unit is complete with extension cards and cables.

Attachment 1: IMG_1364.JPG
IMG_1364.JPG
  6843   Thu Jun 21 02:57:16 2012 JenneBureaucracyGeneral40m Action Items

PRIORITY ITEMS:

ASS doesn't run on Ubuntu!
Put beatbox back, simultaneous arm ALS
Input matricies, output filters to tune SUS.  check after upgrade.
POX11 whitening is not toggling the analog whitening???
Look into PMC PZT drift - PZT failing? Real MC length change?
Vent planning / organization

THE FULL LIST:

General

     Audio system for the signals!!!! Even a crappy one!

SUS

     Input matricies, output filters to tune SUS.  Check after upgrade.

IFO

     Fix occasional common-mode power transient in the arm transmissions. Probably an alignment thing. Would ISS help?

     Drift of the green incident axis -> Assess the amount of the drift / replace the mount

     Calibration of POP22 / AS110

     PMC/IMC/ARM characterization (loss, finesse, reflectivity, etc)

     Arm cavity sweeps, mode scan

     Align AS OSA (others?)

     Investigate PRMI glitches, instability

     PZT or Picomotor mounts for PSL/ALS beams

     ALS on the both arm simultaneously / common / diff ALS scripts

     Measure green locking (Aux laser to arm) transfer functions, residual spectra

     Measure oplev spectra while locking Xgreen - see if the optics are particularly noisy

     Measure Xarm residual motion using POX while ALS is engaged.

     Fix Vio2 filter modules on SUS 

     Switch power supply for amplifiers of beatnote signal to rack power

     Add temp sensors for end lasers to CDS slow channels

     Put windows / pickoffs on PSL table for (a) green trans camera, (b) GTRY, (c) GTRX

CDS

     Capture OSA signals in CDS (the 'scope TDS1001B has a USB port in the back for connecting to the computer)

     Transmon (arms) for high and low power

     POX11 whitening is not toggling the analog whitening???

     Install guardians to monitor EPICS values

Electronics

     Actuator noise level characterization (coil driver response in V/m & coil driver noise level V/rtHz)

     Improvement of POP22/110/AS110 RF circuits?

MEDM

     Complete 40m overview screen - everything should be clickable with pseudo 3D icons

     Script to generate a MEDM tree

     Resurrect MEDM snapshots

     New ! buttons on every screen, include wiki page

     Add all screens to svn

SCRIPT

     Daily diagnosis of the MC spot positions (there must be something already...)

     Daily/occasional adjustment of the incident axis on the MC

     Panic button on Watchdog screen isn't working on Ubuntu

     OPLEV/OSEM trending script before the IFO work for diagnosis. Put into 40m summary screen.

     Auto-locker for arms

     Auto-locker for PSL things

     Diagnostic script for CDS - mx_stream, other stuff.

     Make sure scripts are all svn-ed

Video

     If each video screen has a caption, that would be great

     GUI interface of "videoswitch"

Ubuntu vs. CentOS

     Upgrade Ottavia to Ubuntu, make sure connect to DTT, Dataviewer, AWG.

MM

     IPPOS beam measurement

     AS beam measurement (if beam is bright enough)

     Mode matching calculations, sensitivity to MC waist measurement errors, PRM position

     Think up diagnostic measurement to determine mode matching to PRC while chambers are open, while we tweak MMT

     Use sensoray to capture, measure beam mode at AS, POP

Stray Light

     Scattered light measurement at the end stations: design / confirmation of the mechanical parts/optics/cameras

     Align AUX laser into dark port

     Assemble in-vac beam dumps - how many do we need?

OAF

     OAF comparison plot, both online and offline, comparing static, adaptive and static+adaptive

     Static-only OAF noise budget (Adaptive noise budget as next step)

     Script for daily / weekly re-calculation of Wiener, post to elog if need changing

Tip Tilts

     Prepare electronics for TTs (coil drivers)

      In-air TT testing to confirm we can control / move TTs before we vent

     Connect TTs to digital system and controls, lay cables if needed

     Determine whether we need to add a new flange to OMC chamber

RF Photodiodes

     Opto Energy diode laser - purchase

     Set everything up

     Demod board for AS110 - so we can also have POP110?

  6842   Thu Jun 21 01:58:29 2012 JenneUpdateGreen LockingXgreen preparations

[Yuta, Koji, Jenne]

Lots of small things happened tonight, in preparation for having both arms' ALS working simultaneously.

1. Xarm aligned in IR

     1.1 ETMX oplev centered

2. Xgreen coarsely aligned to Xarm

3. X beat setup on PSL table resurrected. 

     3.1 Steering optics for both X and Y green (before PBS) were touched to fix clipping Xgreen on some of the first mirrors after the light exits the chambers.

     3.2 Xgreen aligned to beat PD

     3.3 PSL green waveplate rotated so ~half of the light goes to X beat, other ~half goes to Y beat (recall we had rotated the polarization so we had max light on the Y beat PD a few weeks ago).

          3.3.1 Now we have ~80uW of PSL green going to each beat PD.

     3.4 PSL green aligned to X beat PD

          3.4.1 Replaced mount for mirror between PBS (which splits PSL green light) and BS (which combines PSL green and X green) so that I could get the alignment correct without having to use the full range of the knobs on the mount.

     3.5 Realigned (coarsely) Ygreen to Y beat PD - the mirrors just after the chambers had been touched, so Y green was no longer directly on the PD.  This will need to be done more finely when we're ready to lock the Yarm again.

     3.6 Dedicated cables for the DC of each beat PD were put in place, so we have those in addition to the DC transmission PDs which we are putting in temporarily each time we align the green to the cavities.  Some mystery unused cables that were running under the PSL table were removed.  The power for the X beat PD was rerouted so that it's much closer to the actual diode, and out of the way. 

4. Better alignment of X green to X arm.

     4.1 Put Green Transmission camera into place

     4.2 Noticed that the X green spot on the transmission camera is not nearly as steady as the Y green.  Increased the gain of the X green refl PD on the end table to see if it helped the spot be more steady, but it's still very wiggly.  We reverted the gain to what it was.  We need to fix this!!!!

    4.3 Removed camera, looked at X transmission DC (PD is temporarily in front of the beat PD), tried to increase the transmission.

     4.4 Aligning the green to the X arm has been really tough - there were a few more iterations of camera then DC PD.

     4.5 Measured X green power on the PSL table - 02 mode was ~150uW.  The 00 mode is still not very stable, which is frustrating, although we have a reasonable amount of power transmitted.   

     4.6 The X end green shutter was moved out of the beam path since the green beam was clipping while going through the shutter.  We need to put it back now that the beam is pretty much aligned.  The beam size and the aperture are roughly the same, so we should look to see if there is a different place on the table where the beam is a little smaller, where we can put the shutter.

5. Whitening filters (Pomona box-style) made for the Xarm I and Q channels - these are the same as the whitening for the Y arm.

6. 30m SMA cable made to be used for 2nd delay line.

     6.1 Steve reminded me this morning that we returned one of the fancy spools of cable that was purchased for the delay lines, since it was defective.  We didn't get it replaced because there was debate as to what is the best kind of cable to use.  We need to come to a conclusion, but for now we have a regular RG-405 cable.

7.  Jamie has started work on modifying the beatbox so that we can have 2-arm ALS.  Hopefully that will be done soon-ish, because we're otherwise pretty close to being ready.

  6841   Wed Jun 20 18:43:57 2012 yutaUpdateLSCcalibrated POX error signal

[Jenne, Yuta]

We did the same calibration for POX. It was 3.8e12 counts/m. See elog #6834 for the details of calibration we did.

According to Kiwamu's calibration, actuator response of ITMX is;

A_ITMX  = 4.913e-09 Hz^2*counts/m / freq^2

Plots below are results from our calibration measurement.

LSCxarmTF_usingITMX.pngLSCxarm_HAover1plusG.pngPOXerrorcalibration.png

  6840   Wed Jun 20 18:09:23 2012 yutaUpdateLockingboth arms aligned, ITMX oplev centered

[Jenne, Yuta]

We aligned FPMI. I also centered ITMX oplev because the light was not hitting on QPD.
Alignment procedure we took was;

1. Align Y arm to the Y end green(Y green trans to PSL is now 195 uW with Y end laser measured temperature 34.14 degC).
2. Aligned IR using PZT2 to Yarm(Now, TRY ~ 0.90).
3. Aligned ITMX monitoring AS spots.
4. Aligned X arm so that TRX maximize.
5. Fine adjusted both BS and X arm(Now, TRX ~ 0.82).

Beam spot position on ETMX looks a little too high & left (from ETMXF camera), but we will leave it until ASS scripts is fixed.

FPMIalignment2010620.png

  6839   Wed Jun 20 18:00:10 2012 ranaSummaryPSLsummaries

 Nice PSL summaries from LHO:

https://alog.ligo-wa.caltech.edu/aLOG/index.php?callRep=3187

  6838   Wed Jun 20 16:37:11 2012 yutaUpdateLockingETMX 1064 trans camera

[Jenne, Yuta]

We made ETMXT camera working.
We connected the camera to video mux, placed 10% pick off mirror in front of TRX PD, lead the beam go to ETMXT camera.
Transmission to the TRY PD was 23.8 uW, but now, it's 21.3 uW (2.3 uW goes to the camera).
So, we changed C1:LSC-TRX_GAIN from -0.00181818 to -0.00203158 (=-0.00181818*23.8/21.3).

There is a channel for power normalization, C1:LSC-TRX_POW_NORM, but is 1 and it looks like we are using this gain for the normalization. Situation of TRY is the same as TRX.

  6837   Wed Jun 20 01:02:20 2012 JamieUpdateGreen LockingRF amp removed from X arm ALS setup

I very badly forgot to log about this in the crush of surfs.

I removed Koji's proto-beatbox RF comparator amp from the X arm ALS setup.  I was investigating hacking it onto one of the discriminator channels in the new beatbox, now that Yuta/Koji's Yuta/Koji's phase tracker is making the coarse beatbox path obsolete.  Upon further reflection we decided to just go ahead and stuff the beatbox board for the X arm, and use the proto-beatbox to test some faster ECL comparators.  This will be done first thing in the morning.

In the meantime the old amp is in my cymac mess on the far left of the electronics bench.

  6836   Wed Jun 20 00:02:16 2012 yutaUpdateGreen Lockingslower scan using phase tracking ALS

For those of you who want to see plots from slower scan.

YarmScan20120619.png

  6835   Wed Jun 20 00:01:04 2012 JenneUpdateLSCcalibrated POY error signal

[Yuta, Jenne]

We have measured the out of loop residual motion of the Yarm while locked with the ALS.  We see ~70pm RMS, as compared to Kiwamu's best of ~24pm RMS.  So we're not yet meeting Kiwamu's best measurement, but we're certainly not in crazy-land.

The Yarm ALS was locked, I took a spectrum of POY11_I_ERR, and used the calibration that we determined earlier this evening.  For reference, I attach a screenshot of our ALS loop filters - we had on all the boosts, and both resonant gain filters (~3Hz and ~16Hz).

A large part of the RMS is coming from the 60Hz power line and the 180Hz harmonic....if we could get rid of these (how were they eliminated from the measurement that Kiwamu used in the paper?? - plotted elog 6780) we would be closer. 

Also, it looks like the hump (in our measurementf ~100Hz, in Kiwamu's ~200Hz) is not quite an order of magnitude higher in amplitude in our measurement vs. Kiwamu's.  We have ~5e-11 m/rtHz, Kiwamu had ~7e-12 m/rtHz.  This increase in noise could be coming from the fact that Yuta and Koji decreased the gain in the Ygreen PDH loop to prevent the PDH box from oscillating. 

While we should still think about why we can't use the same gain that Kiwamu was able to ~6 months ago, we think that we're good enough that we can move on to doing mode scans and residual motion measurements of the Xarm.

 

Attachment 1: LSC_POY_11_I_ERR_calib_19June2012.pdf
LSC_POY_11_I_ERR_calib_19June2012.pdf
Attachment 2: POY_calib_19June2012_FiltBankSettings.png
POY_calib_19June2012_FiltBankSettings.png
  6834   Tue Jun 19 23:36:19 2012 yutaUpdateLSCcalibrated POY error signal

[Jenne, Yuta]

We calibrated POY error signal(C1:LSC-POY11_I_ERR). It was 1.4e12 counts/m.

Modeling of Y arm lock:
  Let's say H is transfer function from Y arm length displacement to POY error signal. This is what we want to measure.
  F is the servo filter (filter module C1:LSC-YARM).
  A is the actuator TF using ITMY. According to Kiwamu's calibration using MICH (see elog #5583),

  A_ITMY  = 4.832e-09 Hz^2*counts/m / freq^2

  We used ITMY to lock Y arm because ITMY is already calibrated.

What we did:
  1. Measured openloop transfer function of Y arm lock using POY error signal using ITMY (G=HFA). We noticed some discrepancy in phase with our model. If we include 1800 usec delay, phase fits well with the measurement. I think this is too big.
LSCyarmTF_usingITMY.png


  2. Measured a transfer function between actuator to POY error signal during lock. This should give us HA/(1+G).
LSCyarm_HAover1plusG.png

  4. Calculated H using measurements above. Assuming there's no frequency dependance in H, we got

  H = 1.4e12 counts/m

POYerrorcalibration.png

 For sanity check; Peak to peak of the POY error signal when crossing the IR resonance is about 800 counts. FWHM is about 1 nm, so our measurement is not so crazy.

  6833   Tue Jun 19 20:26:50 2012 JenneHowToLockingSummer Plan

Jenne and Yuta's Summer Plan

These are the things that we'd like to accomplish, hopefully before Yuta leaves in mid-July

* Yarm mode scan

  ~ Measure residual motion of Yarm cavity when ALS is engaged

* Xarm mode scan

  ~ Align Xarm IR

  ~ Align Xarm green to cavity

  ~ Do mode scan (similar to Yarm)

  ~ Measure residual motion of Xarm cavity when ALS is engaged

* Hold both arms on IR resonance simultaneously (quick proof that we can)

  ~ Modify beatbox so we can use both X and Y at the same time (Jamie will do this Wednesday morning - we've already discussed)

* PRMI + Arms

  ~ Lock the PRMI (which we already know we can do) holding arms off resonance, bring both arms into resonance using ALS

* PRC mode matching - figure out what the deal is

  ~ Look at POP camera with video capture - use software that Eric the Tall wrote with JoeB to measure spot size

* DRMI glitches

  ~ Why can't we keep the DRMI locked stably?

* DRMI + Arms

  ~ Full lock!!

  ~ Make lots of useful diagnostics for aLIGO, measure sensing matricies, etc.

  6832   Mon Jun 18 23:54:31 2012 yutaUpdateGreen Lockingphase tracker for ALS

[Koji, Jenne, Yuta]

Summary:
  We put phase tracker in FINE loop for ALS. We checked it works, and we scanned Y arm by sweeping the phase of the I-Q rotator.
  From the 8 FSR scan using FINE (30 m delay line), we derived that Y arm finesse is 421 +/- 6.

What we did:
  1. We made new phase rotator because current cdsWfsPhase in CDS_PARTS doesn't have phase input. We want to control phase. New phase rotator currently lives in /opt/rtcds/userapps/trunk/isc/c1/models/PHASEROT.mdl. I checked that this works by sweeping the phase input and monitoring the IQ outputs.

  2. We made a phase tracker (/opt/rtcds/userapps/trunk/isc/c1/models/IQLOCK.mdl) and included in c1gcv model. Unit delay is for making a feed back inside the digital system. Currently it is used only for BEATY_FINE (Simulink diagram below). We edited MEDM screens a little accordingly.
newIQLOCK.png


  3. Phase tracking loop has UGF ~ 1.2 kHz, phase margin ~50 deg. They are enough becuase ALS loop has UGF ~ 100 Hz. To control phase tracking loop, use filter module C1:ALS-BEATY_FINE_PHASE (with gain 100). Sometimes, phase tracking loop has large offset because of the integrator and freedom of 360*n in the loop. To relief this, use "CLEAR HISTORY."

  4. Locked Y arm using C1:ALS-BEATY_FINE_PHASE_OUT as an error signal. It worked perfectly and UGF was ~ 90 Hz with gain -8 in C1:ALS-YARM filter module.

  5. Swept phase input to the new phase rotator using excitation point in filter module C1:ALS-BEATY_FINE_OFFSET. Below is the result from this scan. As you can see, we are able to scan for more than the linear range of FINE_I_IN1 signal. We need this extra OFFSET module for scanning because BEATY_FINE_I_ERR stays 0 in the phase tracking loop, and also,  error signal for ALS, output of PHASE module, stays 0 in ALS loop.
YarmScan20120618.png

  6. We analyzed the data from 8FSR scan by FINE with phase tracker using analyzemodescan.py (below). We got Y arm finesse to be 421 +/- 6 (error in 1 sigma). I think the error for the finesse measurement improved because we could done more linear sweep using phase tracker.
fine8FSRscan.png


Next things to do:
  - Phase tracker works amazingly. Maybe we don't need COARSE any more.
  - Install it to X arm and do ALS for both arms.
  - From the series of mode scan we did, mode matching to the arm is OK. There must be something wrong in the PRC, not the input beam. Look into PRC mode matching using video capture and measuring beam size.

  6831   Mon Jun 18 23:38:39 2012 JenneUpdateLSCAdded LSC channels to frames

Since the .ini files get overwritten every time a model is compiled now, we need to put all channels we want saved to frames in the DAQ Channels list inside the model.

I added the _ERR channels for all RFPDs (I and Q for each), as well as the _OUT channels for the DCPDs.  I also added the _OUT channels for the DoF servos (ex. C1:LSC-DARM_OUT).  I don't remember off the top of my head what else we used to save from the LSC model, but those all seemed like ones we'll possibly want access to later. 

We need to go through and do this to all the models we use regularly.

Since SUS hasn't been recompiled in a while, all those channels are saved (until such time as someone does a recompile).  Den has gone through and edited the PEM and OAF .ini files by hand each time he recompiles, so we have that data, although we need to put it into the model (which is the new proper way to acquire channels).

  6830   Mon Jun 18 17:28:03 2012 yutaSummaryComputersbugs in CDS_PARTS/simLinkParts/Fcn

Fcn module in CDS_PARTS is used to include a user defined function in a model.
We should be able to use this by entering desired function, but I found some bugs.

BUG1: Fcn doen't work without ";"

If you put ";" after the function, we can compile.

 sin(u[1]);

But if you put without ";", like

 sin(u[1])

you get the following error message when compiling.

controls@c1ioo
~ 0$ rtcds make c1gcv
### building c1gcv...
Cleaning c1gcv...
Done
Parsing the model c1gcv...
Done
Building EPICS sequencers...
Done
Building front-end Linux kernel module c1gcv...
echo >> target/c1gcvepics/README.making_changes
echo 'Built on date' `date` >> target/c1gcvepics/README.making_changes
make[1]: Leaving directory `/opt/rtcds/caltech/c1/rtbuild'

make[1]: Entering directory `/opt/rtcds/caltech/c1/rtbuild/src/fe/c1gcv'
make -C /lib/modules/2.6.34.1/build SUBDIRS=/opt/rtcds/caltech/c1/rtbuild/src/fe/c1gcv modules
make[2]: Entering directory `/usr/src/linux-2.6.34.1-cs'
  CC [M]  /opt/rtcds/caltech/c1/rtbuild/src/fe/c1gcv/c1gcv.o
make[2]: Leaving directory `/usr/src/linux-2.6.34.1-cs'
make[1]: Leaving directory `/opt/rtcds/caltech/c1/rtbuild/src/fe/c1gcv'
/opt/rtcds/caltech/c1/rtbuild/src/fe/c1gcv/c1gcv.c: In function 'feCode':
/opt/rtcds/caltech/c1/rtbuild/src/fe/c1gcv/c1gcv.c:615: error: expected expression before ';' token
make[3]: *** [/opt/rtcds/caltech/c1/rtbuild/src/fe/c1gcv/c1gcv.o] Error 1
make[2]: *** [_module_/opt/rtcds/caltech/c1/rtbuild/src/fe/c1gcv] Error 2
make[1]: *** [default] Error 2
make: *** [c1gcv] Error 1


BUG2: sindeg doesn't work properly

sindeg should work as cosine with input in degrees.
I made a simple model to test this(below).
model_sindegbug.png


Output of the filter module C1:ALS-BEATY_FINE_PHASE goes to "PHASE_in"
sindeg of this goes to C1:ALS-BEATY_FINE_I_ERR
cosdeg of this goes to C1:ALS-BEATY_FINE_Q_ERR

If you sweep the phase input, you should get sin and cos, but you get the following.
cosdeg (C1:ALS-BEATY_FINE_Q_ERR) looks OK, but sindeg (C1:ALS-BEATY_FINE_I_ERR) looks funny. It looks like ~20000 is its period.

dv_sindegbug.png

  6829   Mon Jun 18 16:23:59 2012 JenneUpdateLockingLSC trigger update

The LSC triggers for the individual filter modules in a filter bank now works.  This is handy so that boosts can come on as soon as a cavity is locked, but will turn off when the cavity unlocks.

You choose which filter modules you want to be triggered, and which ones you want to be manually controlled. 

Example:  LSC-YARM    FM4 and FM5 should always be on, but FM2 and FM3 are controlled by the trigger.  You can set the trigger thresholds for the filter modules independently of the main DoF enable trigger thresholds.

  6828   Mon Jun 18 02:31:43 2012 yutaSummaryGreen Lockinganalysis of mode scan data

I analyzed mode scan data from last week.
Mode matching ratio for Y arm is 86.7 +/- 0.3 %. Assuming we can get rid of TEM01/10 by alignment, this can be improved up to ~ 90%.

Peak search, peak fitting and finnesse calculation:
  I made a python script for doing this. It currently lives in /users/yuta/scripts/modescanresults/analyzemodescan.py.
  What it does is as follows

  1. Read mode scan data(coarse5FSRscan.csv, fine1FSRscan.csv). Each column in the data file should be

[time] [some thing like C1:ALS-BEAT(Y|X)_(COARSE|FINE)_(I|Q)_IN1] [C1:LSC-POY11_I_ERR] [C1:LSC-TRY_OUT]

Each separated by comma. Currently, this script uses only TRY, but it reads all anyway

  2. Find peak in TRY data. For the peak search, it splits data in 1 sec and find local maximum. If the local maximum is higher than given threshold, it recognize it as a peak. If two peaks are very close, it uses higher one. This sometimes fails, because mode scan data we have is not so nice.

  3. Fit each peak with Lorentzian function,

TRY = a*b/(4*(t-c)^2+b^2) + d  (a>0, b>0)

  where a/b is a peak height, b is a linewidth (FWHM), c is a peak position in time, and d is a offset.
  I don't like this, but currently, a/b+c is fixed to the maximum value of TRY data used for fitting. This is because sometimes TRY data is so bad and I couldn't get the peak height correctly. Each points of TRY data doesn't have same error because cavity length is fluctuating and relation between cavity length and TRY is not linear. I think I should use some weighting for the fit, but currently, I just use least squares.

  4. Find TEM00 and calculate FSR in "seconds". I just used "seconds" assuming we did a linear sweep. This script recognize TEM00 from the given threshold.

  5. Calculate finesse using FSR and linewidth of the closest TEM00.

  Below are the result plots from this analysis. Calculated finesse looks quite high (~1000). I think this is from non-linearity in the sweep and error in "measured" line width.
coarse5FSRscan.pngfine1FSRscan.png


Higher order modes and RF sidebands:

  Assuming the curvature of ITMY/ETMY are flat/57.5 m, Y arm length is 38.6 m(FSR 3.9 MHz), positions of HOMs and RF sidebands(11/55 MHz) in frequency domain should look like the plot below.
  The script for calculating this currently lives in /users/yuta/scripts/modescanresults/HOMRFSB.py, inspired by Yoichi's script for KAGRA
HOMRFSB.png

Mode-matching ratio:
  By comparing mode scan data and HOM/RF SB positions in a sophisticated way, you can tell which peak is which.
coarse5FSRscanHOMRFSB.png


  From COARSE 5FSR measurement, peak heights are

TEM00 0.884, 0.896, 0.917, 0.905, 0.911
TEM01 0.040, 0.037, 0.051, 0.054, 0.062
TEM02 0.083, 0.078, 0.079, 0.071, 0.078
TEM03 0.018, 0.015, 0.013, 0.015, 0.014

  So the mode-matching ratio is

MMR = 86.2 %, 87.3 %, 86.5 %, 86.6 %, 85.5 %

  From FINE 1FSR measurement, peak heights and mode matching ratio is

TEM00 0.921
TEM01 0.031
TEM02 0.078
TEM03 0.014

MMR = 88.2 %

  Assuming each measurement had same error, mode-matching ratio from these 6 values is

MMR = 86.7 +/- 0.3 %  (error in 1 sigma)

  This can be improved by ~5% by alignment because we still see ~5% of TEM01/10. Study in systematic errors on going.

  6827   Sat Jun 16 19:32:11 2012 yutaUpdateGreen LockingY arm length using 5FSR scan

I know!
But I think there's some error (~ 10% ?) in calibrating the beatbox. In elog #6815, slope near zero crossing point is about 68 counts/MHz, but now, its 60 counts/MHz. Also, zero crossing point in elog #6815 was 47 MHz, but now, its 45 MHz. 5FSR scan was done between these two calibration measurement.

Quote:

Quote:

Calibrating beat frequency to Y arm length change;
  I used L = 32.36 m (figure above, bottom plot).
    dnu_g / dL = c / lamb_g / L = 1.74 MHz/m

Wow. This is way too short.

You don't need to use Albertoo's arm length as his measurement was done before the upgrade.

 

  6826   Sat Jun 16 18:51:44 2012 KojiUpdateGreen LockingY arm length using 5FSR scan

Quote:

Calibrating beat frequency to Y arm length change;
  I used L = 32.36 m (figure above, bottom plot).
    dnu_g / dL = c / lamb_g / L = 1.74 MHz/m

Wow. This is way too short.

You don't need to use Albertoo's arm length as his measurement was done before the upgrade.

  6825   Sat Jun 16 18:17:00 2012 yutaUpdateGreen LockingY arm length using 5FSR scan

Calibrating error signal to beat frequency;
  I injected 0 dBm RF sine wave into the beatbox and sweeped the frequency(just like we did in elog #6815).
  This time, we have different whitening filters. I sweeped the frequency from 0 to 100 MHz in 200 sec.
  The length of the delay line is ~1.5 m for COARSE.
ALS-BEATY_COARSE_I_IN1_DQ.png

Y arm length;
  Here, I think we need some assumption. Let's assume wavelength of IR lamb_IR = 1064 nm and Y end green frequency is nu_g = 2*nu_IR.
  There is a relation
    dnu_g / nu_g = dL / L
  So,
    dnu_g / (dL/lamb_IR) = 2*nu_IR * lamb_IR / L = 2c/L
  We know that dL/lamb_IR = 1/2 for difference in beat frequency between TEM00s. Therefore, slope of the dnu_g vs dL/lamb_IR plot gives us the arm length L(figure below, middle plot).

CalibYarmScan20120614_2.png

  Error estimation is not done yet, but I think the COARSE_I_IN1 error signal to the beat frequency calibration has the largest error because it seems like the amplitude of sine wave changes ~10% day by day.

Calibrating beat frequency to Y arm length change;
  I used L = 32.36 m (figure above, bottom plot).
    dnu_g / dL = c / lamb_g / L = 1.74 MHz/m

  6824   Sat Jun 16 13:01:17 2012 yutaUpdateGreen Lockingscanned Y arm for 5FSR

Quote:

Is that time stamp really correct?

 Yes. I used pyNDS to get data, but here's a screenshot of dataviewer playing back 300 seconds from GPS time 1023780144.


YarmScanDV.png

  6823   Sat Jun 16 12:03:41 2012 ZachUpdateGreen Lockingscanned Y arm for 5FSR

Is that time stamp really correct? I wanted to look at the signal closely to see if I could get any feeling for why it would look so different when positive vs. negative, but I do not see a triangle anywhere near this time (1023780144)...

Quote:

Interesting. It seems for me that there is a dependence of the noisiness as the beat frequency is scanned.

As you increase (or decrease?) the offset, C1:ALS-BEATY-COARSE_I_IN1 becomes bigger and more crisp.
The resulting out-of-loop stability also seems to be improved as you can see from the crispness of the PDH signal.

Do you find why this happens? Is this because the beat S/N depends on the beat frequency due to the PD noise?

 

 

  6822   Sat Jun 16 01:03:21 2012 yutaUpdateGreen Lockingused longer delay line for mode scan

[Mengyao, Yuta]

Last night, I used 1.5 m delay line COARSE and got 5FSR mode scan. The range 5FSR was limited by the range of SR560.
So, this time, we used 6.4 m(21 feet) cable as a delay line for FINE servo. This should increase the sensitivity by factor of 4. But the range will be 4 tmes smaller, ~ 1.3FSR.

Below is the plot of the mode scan.
You can see the peak height difference between TEM00s, but it's just from the resolution of pixels.

You still can see noisiness goes up when blue plot goes down. But this time, 2000 stands for 27 MHz and -2000 stands for 15 MHz in the beat frequency because we flipped the filter gain this time.
Last night, the top of the triangle was about 40 MHz and bottom was about 60 MHz.


YarmScan20120615.png

We are going to derive mode-matching and some cavity parameters using this plot.

  6821   Fri Jun 15 13:33:39 2012 yutaUpdateGreen LockingADC noise contribution to ALS

ADC noise is not a limiting noise source in a current ALS setup.

Below is the calibrated spectrum of C1:ALS-COARSE_I_ERR when
  Y arm swinging with just damping (red; taken last night)
  terminated before AA (green)
  blocked PSL green beam (blue)

Blue and green curve tells us that noise from the beat PD to ADC is not contributing to the Y arm length sensing noise.

YarmALSnoise20120615.png

  6820   Fri Jun 15 01:53:05 2012 KojiUpdateGreen Lockingscanned Y arm for 5FSR

Interesting. It seems for me that there is a dependence of the noisiness as the beat frequency is scanned.

As you increase (or decrease?) the offset, C1:ALS-BEATY-COARSE_I_IN1 becomes bigger and more crisp.
The resulting out-of-loop stability also seems to be improved as you can see from the crispness of the PDH signal.

Do you find why this happens? Is this because the beat S/N depends on the beat frequency due to the PD noise?

 

  6819   Fri Jun 15 00:50:54 2012 yutaUpdateGreen Lockingscanned Y arm for 5FSR

I scanned Y arm for 5FSR (below).
I could done this after I put a whitening filter.
Currently, whitening filter between the beatbox and AA filter is made of

  Ponoma blue box(passive filter with zero at 1 Hz, pole at 10 Hz) + SR560(flat gain 100)

I couldn't do more than 5FSR because SR560 overloads. I checked it by staring at the indicator during the scan.
Reason why we kept loosing lock last night was the overload of  SR560. Mystery solved!

Anyway, 5FSR is enough.
Our next step is to reduce residual arm length fluctuation.

YarmScan20120614_2.png


Also, I increased the alingnment of IR. So, the higher order modes are less than the last scan.

  6818   Thu Jun 14 21:37:37 2012 yutaUpdateGreen Lockingsucceeded in 1FSR mode scan

[Jenne, Yuta]

We couldn't scan the Y arm for 1FSR last night because the ALS servo breaks while sweeping.
We thought this might be from the amplitude fluctuation of the beat signal. The amplitude of the beat signal goes into the beatbox was about -5 dBm, which is not so enough for the beatbox to get good LO. So, we put an amplifier (and attenuators) and the amplitude became +1 dBm. The range beatbox can handle is about -3 dBm to +3 dBm, according to our calculation.

This increased stability of the lock, and we could scan the arm for 1FSR. Below is the plot of scanned ALS error signal (blue), Y arm IR PDH signal (green) and TRY (red).

YarmScan20120614.png

For each slope, we can see two TEM00 peaks, some higer order modes(may be 01, 02, 02) and sidebands (large 11MHz, small 55MHz?).

We couldn't scan for more. This is still a mystery.

Also, we need to reduce residual Y arm length fluctuation more because we get funny TRY peak shape.

Scan speed:
  For C1:ALS-BEATY_COARSE_I_IN1, 1 count stands for 0.21 nm(see elog #6817). We sweeped 4000 peak to peak in 50 sec. So, the scan speed is about 17 nm/sec.
  This means it takes about 0.06 sec to cross resonant peak.
  Cavity build up time is about 2LF/(pi*c) ~ 40 usec. So, the scan is quasi-static enough.
  Characteristic time scale for the Y end temperature control is about 10 sec, so Y end frequency is following the Y arm length change with temperature control.

  Currently, sampling frequency of DQ channels are 2048 Hz. This means we have 100 points in a TRY peak. I think this is enough to get a peak height.

Next step:
  - Reduce RMS. We are trying to use a whitening filter.
  - Find why we can't scan more. Why??
  - ETMY coil gains may have some unbalance. We need to check
  - Characterize Y end green frequency control. Koji and I changed them last week (see elog #6776).
  - Calculate positions of RF SBs and HOMs and compare with this result.

  6817   Thu Jun 14 04:53:39 2012 yutaSummaryGreen Lockingdesigning ALS loop for mode scan

[[Requirement]]
 Arm cavity FWHM for IR is

  FWHM = FSR / F = c/(2LF) = 8 kHz.

 In cavity length, this is

  L/f * FWHM = 40m/(c/1064nm) = 1.2 nm

 So, to do mode scan nicely, arm length fluctuation during resonant peak crossing should be much less than 1.2 nm.


[[Diagram]]
 Let's consider only ADC noise and seismic noise.
ALSloop.png

* S: conversion from Y arm length to the beat frequency

  dL/L = df/f

 So,

  S = df/dL = f/L = c/532nm/40m = 1.4e7 MHz/m


* W: whitening filter

 We set it to flat gain 50. So,

  W = 50


* D: AD conversion of voltage to counts

 D = 2^16counts/20V = 3300 counts/V


* B: frequency to voltage conversion of the beatbox.

 We measured BWD(elog #6815). When we measured this, W was 10. So, the calibration factor at 0 crossing point(~ 50 MHz) is

  B = 1400*0.048/10/D = 0.0021 V/MHz


* A: actuator transferfunction

 I didn't measure this, but this should look like a simple pendulum with ~ 1 Hz resonant frequency.


* n_ADC: ADC noise

 ADC noise is about

  n_ADC = sqrt(2*LSB^2*Ts) = sqrt(2*(20V/2^14)**2*1/64KHz) = 1.6 uV/rtHz


* n_seis: seismic noise

 We measured this by measuring C1:ALS-BEATY_COARSE_I_IN1. This is actually measuring

  D(WBSn_seis + n_ADC)

 Calibrated plot is the red spectrum below.


* F: servo filter (basically C1:ALS-YARM)

 We need to design this. Stabilized arm length fluctuation is

  x_stab = 1/(1+G)*n_seis + G/(1+G)*n_ADC/(WBS)

 where openloop transferfunction G = SBWDFA.
 Below ~ 50 Hz, n_seis is bigger than n_ADC/(WBS). We don't want to introduce ADC noise to the arm. So, UGF should be around 50 Hz. So, we need phase margin around 50 Hz.
 We also need about 10^3 DC gain to get the first term comparable to the second term.

 Considering these things, openloop transferfunction should look like the below left. Expected error signal when ALS on is the below right. I put some resonant gain to get rid of the peaks which contribute to the RMS (stack at 3.2Hz, bounce at 16.5 Hz).
 Inloop RMS we get is about 0.3 nm, which is only 4 times smaller than FWHM.
ALSopenloop.pngyarmlength.png



[[Discussion]]
 We need to reduce RMS more by factor of ~ 30 to get resolusion 1% of FWHM.
 Most contributing factor to the RMS is power line noise. We might want comb filters, but it's difficult because UGF is at around this region.

 So, I think we need more fancy whitening filters. Currently, we can't increase the gain of the whitening filter because SR560 is almost over loading. Whitening filter with zero at 1 Hz might help.

  6816   Thu Jun 14 01:36:34 2012 yutaUpdateGreen Lockingcan't scan Y arm for 1FSR

[Jenne, Koji, Yuta]

We tried to scan of the Y arm but we couldn't scan for more than 1FSR.
In principle, we can do that because the error signal we are using, C1:ALS-BEATY_COARSE_I_IN1, has the range of ~ 40 MHz, which is about 10FSR (see elog http://nodus.ligo.caltech.edu:8080/40m/6815).

ALS stays for more than 10 min when we don't do the scan. If we put some offset gradually from C1ALS-OFFSETTER2, the lock breaks.
We monitored PZT output of the Y end laser, C1:GCY-SLOW_SERVO1_IN1, but it stayed in the range when scanning. So, there must be something wrong in the ALS loop.

Current in-loop arm length fluctuation is about 0.1 nm RMS (0.5 counts RMS).
Below is the spectrum of the error signal when the ALS is off(green) and on (pink,red). Below ~ 50 Hz, the measurement of the Y arm length is limited by ADC noise (~ 2uV/rtHz).
BEATY_COARSE_LoopOnOff.png

  6815   Wed Jun 13 17:39:13 2012 yutaUpdateGreen Lockingcalibrating the beatbox

[Jenne, Yuta]

We put 0 dBm sine wave to the RF input of the beatbox and linear-sweeped frequency of the sine wave from 0 to 200 MHz using network analyzer (Aligent 4395A).
(We first tried to use 11 MHz EOM marconi)

Whlile the sweep, we recorded the output of the beatbox, C1:ALS-BEATY_(FINE|COARSE)_(I|Q)_IN1_DQ. We made them DQ channels today. Also, we put gain 10 after the beatbox before ADC for temporal whitening filter using SR560s.

We fitted the signals with sine wave using least squares fit(scipy.optimize.leastsq).
Transision time of the frequency from 200 MHz to 0 Hz can be seen from the discontinuity in the time series. We can convert time to frequency using this and supposing linear sweep of the network analyzer is perfect.

Plots below are time series data of each signal(top) and expansion of the fitted region with x axis calibrated in frequency (bottom).

ALS-BEATY_COARSE_I_IN1_DQ.pngALS-BEATY_COARSE_Q_IN1_DQ.png
ALS-BEATY_FINE_I_IN1_DQ.pngALS-BEATY_FINE_Q_IN1_DQ.png


We got

C1:ALS-BEATY_COARSE_I_IN1_DQ = -1400 sin(0.048 freq + 1.17pi) - 410
C1:ALS-BEATY_COARSE_Q_IN1_DQ = 1900 sin(0.045 freq + 0.80pi) - 95

C1:ALS-BEATY_FINE_I_IN1_DQ = 1400 sin(0.89 freq + 0.74pi) + 15
C1:ALS-BEATY_FINE_Q_IN1_DQ = 1400 sin(0.89 freq + 1.24pi) - 3.4

(freq in MHz)

The delay line length calculated from this fitted value (supposing speed of signal in cable is 0.7c) is;

  D_coarse = 0.7c * 0.048/(2*pi*1MHz) =  1.6 m
  D_fine = 0.7c * 0.89/(2*pi*1MHz) = 30 m

So, the measurement look quite reasonable.

FINE signals looks nice because we have similar response with 0.5pi phase difference.
For COARSE, maybe we need to do the measurement again because the frequency discontinuity may affected the shape of the signal.

  6814   Wed Jun 13 11:19:05 2012 steveUpdateSUSIDC receptacles clamped

The MC_ IDC 64 pin cables from sat. amplifiers to junction-interface-board towards  whitening - dewhitening at the back of rack 1 X 5 are finally  clamped with

All other sus cables of the same kind have the correct short latch arm to lock them in for reliable contact.

Attachment 1: IMG_1341.JPG
IMG_1341.JPG
  6813   Wed Jun 13 11:10:56 2012 ranaUpdateGreen Lockingmy first modescan (sort of)

You can easily calculate whether or not the coarse readout will work by thinking about the scan resolution you need given the ADC dynamic range and the whitening filter that you use.

  6812   Wed Jun 13 03:03:38 2012 yutaUpdateGreen Lockingmy first modescan (sort of)

Linear range df of the delay line technique is about df ~ c/(2D). So, the linear range for the fine signal(delay line length D=30m) is about 5 MHz.
Arm cavity FSR = c/(2L) = 3.7 MHz.
So, I think we need phase shifting to do mode scan for more than 2 FSRs by holding the arm length finely with fine servo.
For the coarse (D=1.5m), the linear range is about 100 MHz, so if we can do mode scan using coarse servo, it is OK.

In any case, I think it is nice to have linear signal with fixed slope even if we don't adjust the phase every time.

Quote:

 That sounds goofy.

With the delay line technique, you can get a linear signal over 50 MHz with no phase shifting. What is with all this I/Q stuff?

 

  6811   Wed Jun 13 02:24:02 2012 ranaUpdateGreen Lockingmy first modescan (sort of)

 That sounds goofy.

With the delay line technique, you can get a linear signal over 50 MHz with no phase shifting. What is with all this I/Q stuff?

  6810   Wed Jun 13 02:11:59 2012 yutaUpdateGreen Lockingmy first modescan (sort of)

I stabilized Y arm length by using only I phase coarse signal from the beat(C1:ALS-BEATY_COARSE_I_ERR).
I sweeped the arm length by injecting 0.05Hz sine wave from C1:ALS_OFFSETTER2_EXC.
Below is the plot of TRY and the error signal(ideally, Y arm length) while the sweep.

modescan20120612_1.png

I couldn't hold the arm length tight, so you can see multiple peaks close to each other.
We need to
  - adjust offsets
  - adjust rotation phase of I-Q mixing
  - adjust servo filters

to hold the length tighter.

Also, I couldn't sweep the Y arm length very much. I need to calibrate, but to do the modescan for many FSRs, we need to
  - introduce automatic phase optimizing system
There were sin/cos function in the CDS_PARTS, so I think we can feedback I_ERR to control rotation phase of I-Q mixing.

  6809   Tue Jun 12 23:18:18 2012 yutaUpdateGreen LockingI-Q signals for the beat

[Mengyao, Yuta]

Yes!! We have I-Q signals for the beat!!

What we did:
  1. Aligned Y arm to the Y end green incident beam. The transmission to the PSL was about 195 uW.

  2. Aligned IR beam to the Y arm by adjusting PZTs and got the transmission, C1:LSC-TRY_OUT ~ 0.86.

  3. Aligned green optics on the PSL table to get the beat signal. The beat was found when;

  PSL laser temperature on display: 31.41 deg C
  C1:PSL-FSS_SLOWDC = 1.43
  Y end laser "T+": 34.049 deg C
  Y end laser "ADJ": 0
  Y end laser measured temperature: 34.14 deg C
  C1:GCY-SLOW_SERVO2_OFFSET = 29950
  Y end slow servo: off (was on)

  4. Connected the beat PD output to the beatbox.

  5. Kicked ETMY position to change the cavity length and while the ringdown, we run pynds to get data. We plotted C1:ALS-BEATY_FINE_I_ERR vs C1:ALS-BEATY_FINE_Q_ERR, and C1:ALS-BEATY_COARSE_I_ERR vs C1:ALS-BEATY_COARSE_Q_ERR (below). We got nice circle as expected.

FINEIQplot20120612.pngCOARSEIQplot20120612.png

Current setup:
  Only AA filers are put between the output of the beatbox and the ADC.

beatysetup20120612.png

  6808   Tue Jun 12 20:35:46 2012 yutaUpdateGreen Lockingc1gcv recompiled

[Jamie, Yuta]

We recompiled c1gcv because the order of the channels were confusing. We found some change in the phase rotation module when we did this.

I did some cabling and checked each signals are actually going to the right channel. I labeled all the cables I know, which go into the AA chasis for ADC1 of c1ioo machine.

Below is the list of the channels. If you know anything about "unknown" channels, please let me know.

Current channel assignments for ADC1 of c1ioo machine:
  Red ones were added today. Green ones existed in the past, but channel assignment were changed.

cable

# on AA chassis name in Simulink channel name

connected
but unknown

J1A    
   
not connected J1B    
   
not connected J2 adc_1_2 C1:ALS-XARM_BEAT_DC
not connected adc_1_3 C1:ALS-YARM_BEAT_DC
connected
but unknown
J3    
   
connected
but unknown
J4    
   
connected
but unknown
J5    
   
connected
but unknown
J6    
   
connected
but unknown
J7    
   
beat Y arm fine I J8A adc_1_14 C1:ALS-BEATY_FINE_I
beat Y arm fine Q adc_1_15 C1:ALS-BEATY_FINE_Q
not connected J8B    
   
connected
but unknown
J9A    
   
not connected J9B    
   
connected
but unknown
J10    
   
connected
but unknown
J11    
   
not connected J12 adc_1_22 C1:ALS-BEATX_COARSE_I
not connected adc_1_23 C1:ALS-BEATX_COARSE_Q
not connected J13 adc_1_24 C1:ALS-BEATX_FINE_I
not connected adc_1_25 C1:ALS-BEATX_FINE_Q
beat Y arm coarse I
J14 adc_1_26 C1:ALS-BEATY_COARSE_I
beat Y arm coarse Q adc_1_27 C1:ALS-BEATY_COARSE_Q
not connected J15 adc_1_28 Broken! Don't use this!!
adc_1_29 (not broken)
not connected J16A adc_1_30 (not broken)
adc_1_31 Broken? Funny signal.
not connected J16B    
   

Memorandum for me:
  Recompiling procedure;

ssh c1ioo

rtcds make c1gcv
rtcds install c1gcv
rtcds start c1gcv

Attachment 1: c1gcv20120612-2.png
c1gcv20120612-2.png
  6807   Tue Jun 12 17:46:09 2012 JenneUpdateComputersrtcds: command found

Quote:

Quote:

We can't compile any changes to the LSC or the GCV models since Jamie's new script / program isn't found.  I don't know where it is (I can't find it either), so I can't do the compiling by hand, or point explicitly to the script.  The old way of compiling models in the wiki is obsolete, and didn't work :(

Sorry about that.  I had modified the path environment that pointed to the rtcds util.  The rtcds util is now in /opt/rtcds/caltech/c1/scripts/rtcds, which is in the path.  Starting a new shell should make it available again.

 Added TRX and TRY and POY11_I_ERR and POX11_I_ERR to the c1lsc.mdl using a new-style DAQ Channels block, recompiled, installed, started the model, all good.  Restarted the daqd on the framebuilder, and everything is green.  I can go back and get recorded data using dataviewer (for the last few minutes since I started fb), so it all looks good.

Note on the new DAQ Channels block:  Put the text block (from CDS_PARTS) at the same level as the channel you want to save, and name it exactly as it is in the model.  The code-generator will add the _DQ for you.  i.e. if you define a channel "TRY_OUT_DQ" in the lsc model, you'll end up with a channel "C1:LSC-TRY_OUT_DQ_DQ".

  6806   Tue Jun 12 17:29:28 2012 DenUpdateCDSdq channels

All PEM and IOO DQ channels disappeared. These channels were commented in C1???.ini files though I've uncommented them a few weeks ago. It happened after these models were rebuild, C1???.ini files also changed. Why?

I added the channels back. mx_stream died on c1sus after I pressed DAQ reload on medm screen. For IOO model it is even worse. After pressing DAQ Reload for C1IOO model DACQ process dies on the FB and IOO machine suspends.

I rebooted IOO, restarted models and fb. Models work now, but there might be an easier way to add channels without rebooting machines and demons.

  6805   Tue Jun 12 17:04:55 2012 steveUpdateSUSPRM oplevs servo is still bad

Quote:

Quote:

Yuta claims he fixed the PRM oplev by centering it the other day, but no one has left it on and watched it for a long while, to make sure it's okay.  We watched it now for ~2 min, and it was good, but we're leaving the oplevs off anyway for the night.  Tomorrow we should restore PRM (it's currently restored), turn on the oplevs, and let it sit to make sure it doesn't go crazy.

 

 PRM oplev servo was turned on with PITgain 0.5  and YAWgain  -0.7

Note: gain settings were PIT  1.0  and  YAW --0.5   on Jun 1, 2012 that I measured Feb 23, 2012

 It is still oscillating. Gains turned down to zero.

Attachment 1: PRMolvstillosc.png
PRMolvstillosc.png
  6804   Tue Jun 12 16:33:32 2012 steveUpdateSUSPRM oplevs servo ON for confirmation

Quote:

Yuta claims he fixed the PRM oplev by centering it the other day, but no one has left it on and watched it for a long while, to make sure it's okay.  We watched it now for ~2 min, and it was good, but we're leaving the oplevs off anyway for the night.  Tomorrow we should restore PRM (it's currently restored), turn on the oplevs, and let it sit to make sure it doesn't go crazy.

 

 PRM oplev servo was turned on with PITgain 0.5  and YAWgain  -0.7

Note: gain settings were PIT  1.0  and  YAW --0.5   on Jun 1, 2012 that I measured Feb 23, 2012

  6803   Tue Jun 12 13:49:32 2012 JamieConfigurationComputer Scripts / Programstconvert

A nicer, better maintained version of tconvert is now supplied by the lalapps package.  It's called lalapps_tconvert.  I installed lalapps on all the workstations and aliased tconvert to point to lalapps_tconvert.

ELOG V3.1.3-